OpenStudy (anonymous):

What is the volume of a pyramid with slant height 25 feet and square base with edges of 14 feet?

OpenStudy (anonymous):

Volume = 1/3 * (height * area of the base)

OpenStudy (anonymous):

Given slant height, so need to solve for the height... a^2 + b^2 = c^2 (14/2)^2 + h^2 = 25^2 7^2 + h^2 = 25^2 49 + h^2 = 625 h^2 = 576 h = 24 V=1/3x24x14^2=1568 cubic feet